Consider the following hypothesis test:

Ho: mu<=12

H1: mu>12

A sample of 20 provided a sample mean of 14 and a sample standard deviation s = 4.52 . Use a = 0.05 .

Step 1: Calculate the test statistic for this problem . Round your answer to four decimal places.

Step 2: What is the p-value for your test?

Step 3: State the conclusion to this problem.
Answer #1

Ans:

1)

Test statistic:

t=(14-12)/(4.52/sqrt(20))

t=1.9788

2)

df=20-1=19

p-value=tdist(1.9788,19,1)=0.0313

3)

As,p-value<0.05,we reject the null hypothesis.